### v1

# SocialJax: An Evaluation Suite for Multi-agent Reinforcement Learning in Sequential Social Dilemmas

Abstract
Social dilemmas pose a significant challenge in the field of multi-agent reinforcement learning (MARL). Melting Pot is an extensive framework designed to evaluate social dilemma environments, providing an evaluation protocol that measures generalization to new social partners across various test scenarios. However, running reinforcement learning algorithms in the official Melting Pot environments demands substantial computational resources. In this paper, we introduce SocialJax, a suite of sequential social dilemma environments implemented in JAX. JAX is a high-performance numerical computing library for Python that enables significant improvements in the operational efficiency of SocialJax on GPUs and TPUs. Our experiments demonstrate that the training pipeline of SocialJax achieves a 50× speedup in real-time performance compared to Melting Pot’s RLlib baselines. Additionally, we validate the effectiveness of baseline algorithms within the SocialJax environments. Finally, we use Schelling diagrams to verify the social dilemma properties of these environments, ensuring they accurately capture the dynamics of social dilemmas. Our project is available at https://sites.google.com/view/socialjax.
Machine Learning, ICML
1 Introduction
Solving sequential social dilemmas remains a pivotal challenge in multi-agent reinforcement learning (MARL). Melting Pot (Leibo et al., 2021) and Melting Pot 2.0 (Agapiou et al., 2022) provide comprehensive environments for testing and evaluating a broad range of research topics, such as social dilemmas, reciprocity, resource sharing, and task partitioning. Conducting experiments on Melting Pot is highly challenging, as it requires a large number of environment time steps for training. Consequently, using Melting Pot for experiments requires either extremely high hardware specifications or a substantial amount of computing time. Melting Pot is designed under the assumption that general-purpose hardware (CPU) is inexpensive, whereas specialized accelerator hardware (GPU/TPU) is costly. Under this assumption, the computational efficiency ratio between CPU and TPU using Melting Pot is approximately 100:1. Although JAX (Bradbury et al., 2018) can provide some acceleration for algorithms in Melting Pot, it cannot be used to speed up environment execution since Melting Pot environments run on CPUs. Even if both Melting Pot environments and learning algorithms were implemented in JAX, data would still need to transmit over the network between machines, because Melting Pot was originally designed around remote inference, where the CPUs running the environment are not co-located with the TPUs that are performing learning.
In Melting Pot (Leibo et al., 2021), A3C (Mnih, 2016) is used to adapt to the assumption that CPU resources are vastly more abundant than TPU/GPU resources. In Melting Pot 2.0 (Agapiou et al., 2022), algorithms that are better suited for abundant CPU resources and limited TPU/GPU resources are trained, including ACB (Espeholt et al., 2018), V-MPO (Song et al., 2019), and OPRE (Vezhnevets et al., 2020). Since both the environments and algorithms in our SocialJax implementation can run on GPUs, we can choose PPO (Schulman et al., 2017) as the baseline algorithm. Compared to the A3C/VTrace/ACB family, PPO does not require as much novel experience per update step. PPO employs a clipped loss function to prevent excessive policy updates, enhancing training stability. The A3C/VTrace/ACB family utilizes asynchronous parallelism to improve data collection efficiency, making it well-suited for distributed architectures. However, this approach increases the variance of the policy distribution, which can negatively impact training stability.
(a) Coins
(b) Commons Harvest: Open
(c) Commons Harvest: Closed
(d) Clean Up
(e) Coop Mining
(f) Territory
Figure 1: The SocialJax suite contains 6 multi-agent reinforcement learning environments designed to evaluate social dilemmas. The screenshots in this figure were taken from a third-person visualization perspective. The agents of all environments are restricted to partial observability of their surroundings through a designated observation window.
In RL research, several approaches have been proposed that leverage JAX (Bradbury et al., 2018) to execute environments on GPUs, thereby accelerating computation. PureJaxRL (Lu et al., 2022) leverages JAX to parallelize environments and executes both the algorithms and the environments on the GPU, significantly improving the efficiency of RL. However, PureJAXRL is limited to single-agent environments and does not support multi-agent scenarios. JaxMARL (Rutherford et al., 2023) extends the efficiency gains of JAX to MARL by providing a suite of JAX-based MARL environments and algorithms. While JaxMARL provides JAX-based environments for MARL, most of these environments focus on traditional cooperative settings.
To the best of our knowledge, our work is the first evaluation framework leveraging JAX for sequential social dilemmas. We build multiple social dilemma environments based on JAX, which were previously CPU-based. Moreover, we show that our work significantly accelerates the speed of environment interactions, providing greater convenience and possibilities for research on social dilemma problems.
There has been significant progress in research on sequential social dilemmas, with many studies (Hughes et al., 2018; McKee et al., 2020; Yang et al., 2020) achieving notable breakthroughs. However, one challenge they face is the lack of well-established evaluation standards. The Melting Pot (Agapiou et al., 2022; Leibo et al., 2021) framework presents some basic algorithms, such as A3C (Mnih, 2016), V-MPO (Song et al., 2019), and OPRE (Vezhnevets et al., 2020), but these algorithms are not open-sourced in the architecture of Melting Pot 2.0. Different testing environments for social dilemmas are being widely used, making it challenging to verify algorithm performance on a unified standard. SocialJax provides a different framework for evaluating algorithms targeting social dilemmas, addressing the issue of inconsistent evaluation standards in this field.
Our contributions are as follows: First, we implement six sequential social dilemma environments based on JAX and evaluate the effectiveness of IPPO (De Witt et al., 2020; Rutherford et al., 2023) under both individual and common reward settings. The environments (Figure 1), originated from Melting Pot, include Coins, Commons Harvest: Open, Commons Harvest: Closed, Clean Up, Territory, and Coop Mining, all of which feature commonly studied mixed incentives structures. Second, we show that training IPPO in SocialJax is at least 50 times faster than Melting Pot 2.0 on a machine with 14 CPUs and 1 Nvidia A100 GPU. Third, using Schelling diagrams, we further analyze the behavior of agents trained with individual and collective rewards to understand the underlying incentive structures.
2 Related Work
Sequential Social Dilemmas Libraries (Vinitsky et al., 2019) introduces a testbed for evaluating social dilemmas. However, its scope is restricted to only two environments, Clean Up and Harvest. OpenSpiel (Lanctot et al., 2019) provides a diverse set of environments, including those where cooperative and competitive dynamics coexist, making them ideal for studying social dilemmas. Melting Pot (Leibo et al., 2021) was the first to establish a standardized benchmark for MARL, specifically designed to evaluate agent performance in sequential social dilemmas. By providing a diverse set of carefully designed multi-agent scenarios, it enables systematic testing of agent behaviors in cooperative, competitive, and mixed-motive settings. Melting Pot 2.0 (Agapiou et al., 2022) introduces a more comprehensive and refined suite of testing environments that capture a wide range of interdependent relationships and incentive structures. These improvements allow researchers to better assess generalization, emergent behaviors, and strategic adaptation in MARL, making Melting Pot 2.0 a more powerful tool for studying sequential social dilemmas.
JAX-based RL environments and Algorithms JAX (Bradbury et al., 2018) is an open-source library developed by Google, designed for high-performance numerical computing and machine learning tasks. Based on the GPU and TPU acceleration capabilities introduced by JAX, along with its support for parallelism and vectorization, JAX can effectively accelerate both environments and algorithms on hardware accelerators (GPUs, TPUs). Functorch (Horace He, 2021) provides JAX-like composable function transforms for PyTorch, making it easier to accelerate RL algorithms on hardware accelerators. (Lu et al., 2022) introduces a parallelized framework leveraging JAX, enabling both environment and model training to run on GPUs.
Jumanji (Bonnet et al., 2023) has developed a diverse range of single-agent environments in JAX, ranging from simple games to NP-hard combinatorial problems. Gymnax (Lange, 2022) provides JAX versions of different environments, including classic control, Bsuite (Osband et al., 2019), MinAtar (Young & Tian, 2019), and a collection of classic RL tasks. Pgx (Koyamada et al., 2024) implements a variety of board game simulation environments using JAX, such as Chess, Shogi, and Go. Brax (Freeman et al., 2021) is a physics engine written in JAX for RL, which re-implements MuJoCo. XLand-MiniGrid (Nikulin et al., 2023) is a grid-world environment and benchmark fully implemented in JAX. Similarly, our work is also demonstrated using grid-world environments. Mava (Mahjoub et al., 2024) and JaxMARL (Rutherford et al., 2023) both provide JAX-based MARL environments and baseline algorithms. Among the JAX environments and algorithms, only the Storm and Coins environments in JaxMARL are related to social dilemmas, while none of them specifically address sequential social dilemmas. We construct multiple environments focused on social dilemma challenges.
3 SocialJax
We introduce a JAX version of sequential social dilemma environments that delivers substantial performance gains through parallel processing of both algorithmic computations and environmental simulations on the GPU. Our work presents an efficient testing framework for sequential social dilemmas, complete with evaluation environments and benchmark baselines. Our SocialJax encompasses a diverse suite of environments (Coins, Commons Harvest: Open, Commons Harvest: Closed, Clean Up, Territory, and Coop Mining), which collectively capture a wide range of sequential social dilemma scenarios, including public good dilemmas and common pool resource problems. We evaluate the environments using IPPO and provided an intuitive demonstration of their social dilemma properties through $Schelling$ $diagrams$ (Perolat et al., 2017; Schelling, 1973).
3.1 Sequential Social Dilemmas
In $N$-player partially observable Markov games (Littman, 1994), agents possess partial observation of the environment. At each state, players choose actions from their respective action sets $\mathcal{A}^{1},\ldots,\mathcal{A}^{N}$. State transitions $\mathcal{T}:\mathcal{S}\times\mathcal{A}^{1}\times\ldots\times\mathcal{A}^{N}\to\Delta(\mathcal{S})$ are determined by the current state and the joint actions taken by all agents, where $\Delta(\mathcal{S})$ represents the set of discrete probability distributions over $\mathcal{S}$. We define the observation space of player $i$ as $O_{i}=\{o^{i}|s\in\mathcal{S},o^{i}=O(s,i)\}$. Each agent has its corresponding individual reward defined as $r^{i}:\mathcal{S}\times\mathcal{A}^{1}\times\dots\times\mathcal{A}^{N}\to\mathbb{R}$. In an $N$-player $sequential$ $social$ $dilemma$, cooperators and defectors each have their own disjoint policies$(\pi_{c}^{1},\dots,\pi_{c}^{l},\dots,\pi_{d}^{1},\dots,\pi_{d}^{m})\in\Pi_{c}^{l}\times\Pi_{d}^{m}$ with $l+m=N$. $l$ is the number of cooperators. $R_{c}(l)$ and $R_{d}(l)$ are denoted as the average payoff for cooperating policies and defecting policies. When we have a sequential social dilemma, if it satisfies the following conditions (Leibo et al., 2017; Hughes et al., 2018):
1. Mutual cooperation is greater than mutual defection:
$R_{c}(N)>R_{d}(0).$
2. Mutual cooperation is greater than exploited by defectors:
$R_{c}(N)>R_{c}(0).$
3. Either the fear property or the greed property, or both:
  • Fear: mutual defection is greater than being exploited:
$R_{d}(i)>R_{c}(i)$ for sufficiently small $i$.
  • Greed: exploiting a cooperator is greater than mutual cooperation:
$R_{d}(i)>R_{c}(i)$ for sufficiently large $i$.
We use $Schelling$ $diagrams$ (Perolat et al., 2017; Schelling, 1973) to depict the payoffs of different numbers of cooperators and defectors. The lines $R_{c}(l+1)$ and $R_{d}(l)$ depict the average reward for an agent choosing to either cooperate or defect, as a function of the number of co-players that are cooperating (see Figure 3).
3.2 Environments of SocialJax
Our environments are mainly derived from Melting Pot 2.0 (Agapiou et al., 2022). The Melting Pot suite of environments is developed using an entity-component system implemented in Lua and configured via Python. This system is built on top of the DMLab2D engine, which is written in C++ and configured using Lua (Beattie et al., 2020). Since the core environment computations are always executed in C++ on the CPUs, and the algorithms run on the TPU/GPU, the overall execution speed is constrained by the efficiency of the CPUs and the data transfer bandwidth between the CPUs and GPUs. To enhance the runtime efficiency of the environments, we need to scale up the number of CPUs, which presents significant challenges to the hardware capabilities. In contrast, the environments and algorithms provided by our SocialJax implementation run entirely on GPUs, leveraging the JAX framework to significantly improve execution efficiency.
Figure 1 shows the layout of our environments. The data format observed by agents in SocialJax environments differ from those in Melting Pot 2.0 (Agapiou et al., 2022). SocialJax environments are similar to MiniGrid (Chevalier-Boisvert et al., 2018), where agents observe a grid matrix, with different objects represented by distinct numerical values. In contrast, Melting Pot 2.0 provides observations to agents in the form of image pixels. Melting Pot 2.0 was originally designed to use $8\times 8$ pixels per cell, resulting in observations of $88\times 8$ 88 pixels in most environments. In this setup, the agents have a partially observable window of $11\times 1$ 11 cells. In the Melting Pot Contest 2023 (Trivedi et al., 2024), all academic participants down-sampled the cells to $1\times 1$, also producing observations of $11\times 1$ 11. The agents in our environments have the same $11\times 1$ 11 field of view, but they are implemented at the grid level.
Coins was first introduced by (Lerer & Peysakhovich, 2017) to investigate how agents behave in situations where they can cooperate with or harm others through their actions, making it a classic example of a social dilemma. Two players collect coins in a shared room, where each coin is assigned to one player’s color. Each coin, upon appearing, has an equal 50% probability of being assigned to the first player’s color or the second player’s color. An agent receives a reward of 1 for collecting any coin, regardless of its color. If one player collects a coin assigned to the other player’s color, the other player receives a reward of -2.
Commons Harvest The core concept of Commons Harvest was first defined by (Janssen et al., 2010). (Perolat et al., 2017) introduced it in multi-agent scenarios. Melting Pot 2.0 (Agapiou et al., 2022) has made it a more comprehensive environment. There are several patches of apples in the room. The players can receive a reward of 1, when they collect one apple. The apples will regrow with probability that is determined by the number of neighborhood apples in radius 2, when the apples are collected by the players. The probability of apple regrowth decreases as the nearby apples diminish. Apples will not regrow, if there are no apples in the neighborhood. Commons Harvest includes two scenarios: Commons Harvest: Closed and Commons Harvest: Open.
In Commons Harvest: Closed , the players’ roles are divided into two groups: one focuses on collecting apples, while the other is tasked with preventing over-harvesting by other players (Perolat et al., 2017). There are two rooms in this setup, each containing multiple patches of apples and a single entry. These rooms can be defended by specific players to prevent others from accessing the apples by zapping them with with a beam, causing them to respawn.
In Commons Harvest: Open, all individuals can harvest apples, but must restrict their actions to not collect the last apple and ensure that the apples can regrow (Perolat et al., 2017).
(a) Coins
(b) Commons Harvest: Open
(c) Commons Harvest: Closed
(d) Clean Up
(e) Coop Mining
(f) Territory
Figure 2: Training curves of Common Reward IPPO and Individual Reward IPPO. Common Reward encourages collective interests, leading to higher overall returns, while Individual Reward primarily drives selfish behavior, often resulting in lower returns.
Clean Up is a multiplayer game where players earn rewards by collecting apples, each worth +1 reward. Apples grow in an orchard and their regrowth depends on the cleanliness of a nearby river. Pollution accumulates in the river at a constant rate and once pollution surpasses a certain threshold, the apple growth rate drops to zero. Players have the option to perform a cleaning action that removes small amounts of pollution from the river. For continuous apple growth, the group must keep river pollution levels consistently low over time.
Territory is a competitive multi-agent game where players paint walls to capture territory and earn rewards. The players claim walls by painting them with their unique color, either by touching or flinging paint. The paint dries after 25 steps, brightens, and rewards are given every subsequent timestep. The more walls that are painted and dried, the greater the reward players can receive. Walls can be zapped twice to permanently destroy them, making them non-claimable and traversable. Players can also zap each other and a player is permanently removed after being hit twice.
Nine players compete to claim resources by painting them in their unique colors. Players spawn in an open area near each other, away from the resource walls. Resource distribution is uneven, with some areas contain more walls than others.
Coop Mining In this environment, two types of ore spawn randomly in empty spaces. Players can extract the ore to get reward. Iron ore (gray) can be mined individually and provides a reward of +1 upon extraction. In contrast, gold ore (yellow) requires coordinated mining by two to four players within a 3-step window, granting a reward of +8 to each participant. When a player begins to mine gold ore, the state of the ore changes to “partially mined” to indicate readiness for another player to assist. Visually, this is represented by a brighter shade of yellow. If no other player cooperates or too many players attempt to mine simultaneously, the ore reverts to its original state, and no reward is granted.
(a) Coins
(b) Commons Harvest: Open
(c) Commons Harvest: Closed
(d) Clean Up
(e) Coop Mining
(f) Territory
Figure 3: Schelling diagrams of the environments, visualizing the relationship between individual incentives and collective outcomes. The dotted line represents the overall average return of cooperators and defectors.
3.3 Algorithms in JAX
Due to IPPO’s widespread adoption and stability (Schulman et al., 2017; De Witt et al., 2020), we use it as our baseline algorithm. IPPO is also adopted by PureJaxRL (Lu et al., 2022) and JaxMARL (Rutherford et al., 2023). We leverage convolutional neural networks to process state information from the SocialJax environments. To encourage prosocial behavior, we introduce common rewards, while individual rewards are used to incentivize selfishness in agents.
Common Rewards refers to a scenario where all agents in a multi-agent system share a single, unified reward signal derived from the environment. This approach ensures that all agents are aligned towards achieving the same objective, promoting collaboration and coordination among them. By sharing a common reward, agents are incentivized to work together and make decisions that benefit the collective, rather than focusing solely on individual gains. This can help prevent conflicts and encourage cooperative behavior.
Individual Rewards indicate that each agent is assigned its own reward, inherently encouraging selfish behavior. This approach allows agents to prioritize maximizing their personal success over collective goals. We examine the performance of selfish agents in social dilemma environments and compare it against their performance under a common reward scheme.
Table 1: Results on environment steps per second across various SocialJax environments under random actions. Different environment configurations are tested at scales from 1 to 4096 JAX Env to illustrate performance gains.
Envs 1 Original Env 1 JAX Env 128 JAX Env 1024 JAX Env 4096 JAX Env
Coins $1.2\times 10^{4}$ $2.0\times 10^{3}$ $2.6\times 10^{5}$ $1.4\times 10^{6}$ $3.4\times 10^{6}$
Harvest: Open $3.7\times 10^{3}$ $1.2\times 10^{3}$ $1.1\times 10^{5}$ $5.0\times 10^{5}$ $7.9\times 10^{5}$
Harvest: Closed $3.7\times 10^{3}$ $1.2\times 10^{3}$ $1.1\times 10^{5}$ $5.0\times 10^{5}$ $7.9\times 10^{5}$
Clean Up $2.7\times 10^{3}$ $1.7\times 10^{3}$ $1.8\times 10^{5}$ $9.9\times 10^{5}$ $2.1\times 10^{6}$
Territory $2.5\times 10^{3}$ $6.6\times 10^{2}$ $8.1\times 10^{4}$ $4.6\times 10^{5}$ $9.5\times 10^{5}$
Coop Mining $3.6\times 10^{3}$ $1.9\times 10^{3}$ $1.9\times 10^{5}$ $1.0\times 10^{6}$ $1.5\times 10^{6}$
4 Experiments
4.1 Setup
To validate our library, we perform thorough tests of both Common Reward IPPO and Individual Reward IPPO algorithms across a diverse range of SocialJax environments. In addition to these experiments, we plot Schelling diagrams to further verify that our environments effectively represent social dilemmas. This evaluation ensures that the environments within SocialJax are suitable for studying sequential social dilemmas.
We assess the performance of our JAX-based environment on GPU and compare its speed with the implementations provided by MeltingPot 2.0 (Agapiou et al., 2022). When taking random actions, the speed of SocialJax is approximately 70-400 times faster than Stable-Baselines3 and 40-60 times faster than RLlib, both open-sourced by MeltingPot 2.0.
4.2 Results
Speed Benchmarks From Table 1, we observe a significant increase in the speed of the training pipeline with the addition of more parallel environments. Since both the environments and algorithms are vectorized using JAX, they are well-suited for GPU acceleration. In Table 1, we tested up to 4096 environments in parallel.
We compare the performance of SocialJax and MeltingPot 2.0 (Agapiou et al., 2022) by measuring steps per second on the same hardware, which consists of an NVIDIA A100 GPU and 14 CPU cores. As shown in Table 1, we evaluate the speed of SocialJax across various environments and environment counts using random actions. We observe a significant speedup when running environments in parallel on a GPU compared to running a single environment. The original environments run on a single CPU, and achieving the parallel efficiency of SocialJax would require scaling to hundreds of CPUs. We also compare the performance of our training pipeline with the training speed of the Stable-Baselines3 and RLlib implementations with TensorFlow, based on the publicly available code from Melting Pot 2.0. Note that the experimental results in Melting Pot 2.0 (Agapiou et al., 2022) cannot be directly compared, as the code related to their reported results has not been released. The open-source versions of the algorithms in Melting Pot 2.0 are based on RLlib and Stable Baselines3. However, the versions of Actor-Critic Baseline (ACB) (Espeholt et al., 2018), V-MPO (Song et al., 2019), and OPRE (Vezhnevets et al., 2020) used in their reports have not been made available as open source.
Additionally, since MeltingPot 2.0 is pixel-based environments while our SocialJax environments are grid-based, it is not feasible to align the environment and neural network settings directly. Therefore, we only perform real-time comparisons for training scenarios where convergence is achieved. For relatively simple environments such as Coins, SocialJax can complete 1e9 timesteps within 3 hours, whereas training with Melting Pot 2.0 using Stable-Baselines3 requires approximately 1,300 hours (on an NVIDIA A100 GPU with 14 CPU cores), training with Melting Pot 2.0 using RLlib takes around 150 hours. Consequently, conducting social dilemma experiments with Melting Pot 2.0 is highly challenging without a large amount of CPU resources. Training Coins in SocialJax is about 50-400 times faster compared to Coins in Melting Pot 2.0. For relatively complex environments such as Clean Up and Territory, SocialJax is approximately 50-140 times faster than Melting Pot 2.0.
Algorithm Performance We compare the performance of common reward IPPO and individual reward IPPO across different environments. In the default individual reward setting, each agent receives a reward based on the joint-action, independent of the payoffs of other agents. The environments use homogeneous reward functions for all agents. This creates incentives for agents to free-ride, or to over exploit common pool resources, because only they themselves benefit from their actions, while the social costs are shared by all.
In contrast, a common reward, or team reward means that all agents share a single reward signal, typically the total payoff of all agents in the environment. The common reward encourages cooperation, as the objective of each agent is to maximize the total reward, making individual agents more inclined to adopt collaborative behaviors. In environments such as Coins (Figure 2(a)), Harvest (Figure 2(b), Figure 2(c)), Clean Up (Figure 2(d)), and Coop mining (Figure 2(e)), a common reward typically performs better than individual rewards, as the agents attempt to maximize the total payoff, removing the incentives to engage in selfish behaviors.
However, using a common reward structure can exacerbate the credit assignment problem. Specifically, if an agent takes an arbitrary action concurrently with a teammate who performs a successful action generating a reward, the agent may mistakenly attribute that reward to its own action. This misattribution hinders accurate learning and policy improvement, leading to suboptimal coordination. We observe this phenomenon explicitly in our Territory environment (see Figure 2(f)), where common reward agents fail to identify effective collaborative policies. Consequently, their overall performance remains inferior to agents trained with individual rewards, reflecting a clearer assignment of credit.
4.3 Social Dilemma Environments Attributes
Schelling Diagram Analysis To validate the social dilemma properties of SocialJax, we depict $Schelling$ $diagrams$ for each environment. The cooperative policies are sampled from the agents that used a common reward, while the defector policies originate from agents trained with independent rewards. We evaluate the environments over 30 episodes and compute the average rewards for the cooperative and defector agents.
Coins (Figure 3(a)), Harvest: Open (Figure 3(b)), Harvest: Closed (Figure 3(c)) exhibit fear: namely that agents would prefer to defect if some of their co-players are defecting. An agent actually does better cooperating if all their co-players are cooperating. In Harvest, we observe that it only takes a single defector agent to have a severely negative impact on the returns of other players. This is because defector agents will typically harvest the last apple in a patch, permanently depleting it. The cooperator policies have learned to leave a greater amount of apples in each patch, which causes a higher regrowth rate, so they generally exhibit a greater tendency to abstain from harvesting in the presence of a defector.
In Clean Up (Figure 3(d)), we instead observe greed, as there is an incentive to defect if a sufficient number of co-players are cooperating. This is due to that fact that cooperators will clean the pollution, whereas the defector will not, but it requires a minimum of three cooperators to clean the river for apples to grow. Coop Mining (Figure 3(e)) primarily exhibits the greed property. In this environment, if a sufficient number of agents choose to cooperate by focusing on collecting gold, other agents can obtain a higher payoff by defecting. This is because the cooperators tend to neglect the ore, allowing the defector to collect more ore.
Due to the difficulties of credit assignment in Territory (Figure 3(f)), IPPO with a common rewards fails to learn effective policies: they perform consistently worse than the policies with individual rewards. We believe that this environment should be a social dilemma, because there are personal incentives that undermine the collective welfare, but more advanced algorithms or training methodologies are required to find effective cooperative policies.
5 Conclusion
The use of hardware acceleration has revolutionized MARL research, enabling researchers to overcome computational limitations and accelerate the iteration of ideas. SocialJax is a library designed to leverage these advancements in the context of social dilemmas. It provides a collection of social dilemma environments and baseline algorithms implemented in JAX, offering user-friendly interfaces combined with the efficiency of hardware acceleration. By achieving remarkable speed-ups over traditional CPU-based implementations, SocialJax significantly reduces the experiment runtime. In addition, its unified codebase consolidates a diverse range of social dilemma environments into a standardized framework, facilitating consistent and efficient experimentation and greatly benefiting future research on social dilemmas.
Impact Statement
We present the SocialJax framework, which utilizes JAX to construct a diverse array of social dilemma environments and their corresponding algorithms. Our work offers efficient and varied environments, complemented by a comprehensive training pipeline, thereby facilitating future research in social dilemmas and advancing the field of multi-agent social dilemma studies. It establishes a robust foundation for advancing future community research on social behavior, enabling deeper exploration and broader interdisciplinary collaboration.